From GlueXWiki
Jump to: navigation, search

September 4, 2014 FDC+CDC meeting


  1. Info needed on commissioning.
  2. DAQ
  3. CDC update
  4. FDC status (Vlad, Lubomir)
  5. Electronics (Fernando)
  6. Engineering
  7. Commissioning Wiki Page (all)
  8. Other

Instructions for Bluejeans meeting connection

FDC meeting ID: 290664653


Participants: Curtis, Naomi (CMU), Eugene, Mike, Bill, Chris, Nick, Simon, Fernando, Vlad, and Lubomir (JLab).


- Curtis pointed out to the information needed to be fill in the commissioning plan for the CDC and FDC in the "Tune photon beam" sections. The info for the FDC has already been linked there: it provides detailed instructions for the shifters how to turn ON/OFF the detector including special instructions for doing this after a new beam alignment.


- Lubomir: There were some problems with the DAQ recently, in some cases there was no connection to the ROCs, in other the TDC crates only (!?) were not giving any data. Beni couldn't start it on Wednesday, Sergey also tried. Surprisingly today, without any change, we were able to start the DAQ again.

CDC update

- CDC collected a lot of data last weekend (there were problems with the FDC crates). Mike is analyzing the data and reports several cards not showing data. On one card, the nominal HV value can't be set.


- Fernando with Nick did extensive studies of the noise from the CAEN HV supplies as induced on the CDC channels (see the document linked above). Tests were done on several cards around the noisy one, C6, in different configurations. Rerouting of the HV wires around the HVBs helped significantly in reducing the correlated noise coming from the HV module. As Fernando demonstrated (using the averaging feature of the scope) the noise that is left is a white noise. Similar conclusion was made using spectral analyzer.

- Fernando: the plan is to add ferrite filters to the HV cables (as suggested by the CAEN people). If this doesn't work we may discuss rearranging the HV wires, but this is some work and also will make the access to the HVBs more difficult. Naomi proposed to shield the HV wires, at least around the cards with the highest noise. Fernando: the noise is <14mV with low frequency, it may not be a problem when running with the beam.

FDC update

- Vlad looked at the results from a run taken a month ago and compared with previous results. We are seeing now several tens of channels that were reported bad before: this is because we were able to reduce the threshold down to ~4mV by requiring timing coincidence between the signals coming from the up/down strips and the wires. Due to the low statistics we can't check about 20 channels at the periphery of the chambers. 9 new problematic channels showed up and Vlad will check them with a scope. The problem of one wire channels is most likely the 3M connector; Chris and Nick will look at it.


- Bill: the chiller is working now with the heat exchanger adjusted to have the temperature of the input water of 13.7-13.9 deg. The system has to be tested also with the LV turned off, will do it next week. Bill suspects bad low level switch, that's why it is showing warning all the time, while one can see the level in the sub-tank is not low. Wesley Moore implemented epics monitoring of the chiller temperature, but so far it doesn't turn off the chiller if it is outside the limits.

Due to the special Hall D Worker Safety Awareness Training in the hall we had to stop the meeting at this point.